### v1

# The algebraic and geometric classification of

Abstract: In this paper, we develop a method to obtain the algebraic classification of noncommutative Jordan algebras from the classification of Jordan algebras of the same dimension. We use this method to obtain the algebraic classification of complex $3$-dimensional noncommutative Jordan algebras. As a byproduct, we obtain the classification of complex $3$-dimensional Kokoris, standard, generic Poisson, and generic Poisson–Jordan algebras; and also complex $4$-dimensional nilpotent Kokoris and standard algebras. In addition, we consider the geometric classification of varieties of cited algebras, that is the description of its irreducible components.
Keywords: noncommutative Jordan algebra, generic Poisson algebra, algebraic classification, geometric classification.
MSC2020: 17A30 (primary); 17A15, 17C55, 17B63, 14L30 (secondary).
Introduction
The algebraic classification (up to isomorphism) of algebras of small dimensions from a certain variety defined by a family of polynomial identities is a classic problem in the theory of non-associative algebras. Another interesting approach to studying algebras of a fixed dimension is to study them from a geometric point of view (that is, to study the degenerations and deformations of these algebras). The results in which the complete information about degenerations of a certain variety is obtained are generally referred to as the geometric classification of the algebras of these varieties. There are many results related to the algebraic and geometric classification of Jordan, Lie, Leibniz, Zinbiel, and other algebras (see, [3, 20, 8, 9, 12, 13, 14, 16, 15, 1] and references in [17, 25, 18]).
Noncommutative Jordan algebras were introduced by Albert in [2]. He noted that the structure theories of alternative and Jordan algebras share so many nice properties that it is natural to conjecture that these algebras are members of a more general class with a similar theory. So he introduced the variety of noncommutative Jordan algebras defined by the Jordan identity and the flexibility identity. Namely, the variety of noncommutative Jordan algebras is defined by the following identities:
$\begin{array}[]{rcl}x(yx)&=&(xy)x,\\ x^{2}(yx)&=&(x^{2}y)x\end{array}$
The class of noncommutative Jordan algebras turned out to be vast: for example, apart from alternative and Jordan algebras, it contains quasi-associative and quasi-alternative algebras, quadratic flexible algebras and all anticommutative algebras. However, the structure theory of this class is far from being nice. Nevertheless, certain progress was made in the study of the structure theory of noncommutative Jordan algebras and superalgebras (see, for example [5, 35, 26, 27, 19, 30, 4, 23, 24, 28, 31, 32, 33, 39, 29]). So, Schafer gave the first structure theory for noncommutative Jordan algebras of characteristic zero in [35]. Kleinfeld and Kokoris proved that a simple, flexible, power-associative algebra of finite dimension over a field of characteristic zero is a noncommutative Jordan algebra [22]. A coordinatization theorem for noncommutative Jordan algebras was obtained in a paper by McCrimmon [26]. Later, McCrimmon shows that the property of being a noncommutative Jordan algebra is preserved under homotopy [27]. Noncommutative Jordan algebras with additional identity $([x,y],y,y)=0$ were studied in papers by Shestakov and Schafer [39, 36]. Strongly prime noncommutative Jordan algebras were studied by Skosyrskiĭ [37, 38]. A connection between noncommutative Jordan algebras and $(-1,-1)$-balanced Freudenthal Kantor triple systems was established by Elduque, Kamiya, and Okubo [10]. Cabrera Serrano, Cabrera García, and Rodríguez Palacios studied the algebra of multiplications of prime and semiprime noncommutative Jordan algebras [4, 5]. Jumaniyozov, Kaygorodov, and Khudoyberdiyev classified complex $4$-dimensional nilpotent noncommutative Jordan algebras [16]. Connections between noncommutative Jordan algebras and conservative algebras are discussed in [29]. There are also some results in noncommutative Jordan superalgebras. Namely, Pozhidaev and Shestakov classified all simple finite-dimensional noncommutative Jordan superalgebras in [32, 33]; Kaygorodov, Popov, and Lopatin studied the structure of simple noncommutative Jordan superalgebras [19]; Popov described representations of simple noncommutative Jordan superalgebras in [30, 31].
The main goal of the present paper is to obtain the algebraic and geometric description of the variety of complex $3$-dimensional noncommutative Jordan algebras. To do so, we first determine all such $3$-dimensional algebra structures, up to isomorphism (what we call the algebraic classification), and then proceed to determine the geometric properties of the corresponding variety, namely its dimension and description of the irreducible components (the geometric classification). As some corollaries, we have the algebraic and geometric classification of complex $3$-dimensional Kokoris and standard algebras.
Our main results are summarized below.
Theorem A1.
There are infinitely many isomorphism classes of complex $3$-dimensional noncommutative Jordan algebras, described explicitly in Theorems 9,10,11 and 12 in terms of $9$ one-parameter families, $1$ two-parameter family and $23$ additional isomorphism classes. There are only $18$ non-isomorphic complex $4$-dimensional nilpotent (non-$2$-step nilpotent) noncommutative Jordan algebras listed in Theorem 13.
Theorem A2.
There are infinitely many isomorphism classes of complex $3$-dimensional Kokoris algebras, described explicitly in Theorem 14 in terms of $3$ one-parameter families and $19$ additional isomorphism classes. There are only $10$ non-isomorphic complex $4$-dimensional nilpotent (non-$2$-step nilpotent) Kokoris algebras listed in Theorem 16.
Theorem A3.
There are infinitely many isomorphism classes of complex $3$-dimensional standard algebras, described explicitly in Theorem 17 in terms of $1$ one-parameter family and $32$ additional isomorphism classes. There are only $18$ non-isomorphic complex $4$-dimensional nilpotent (non-$2$-step nilpotent) standard algebras listed in Theorem 13.
From the geometric point of view, in many cases, the irreducible components of the variety are determined by the rigid algebras, i.e., algebras whose orbit closure is an irreducible component. It is worth mentioning that this is not always the case and already in [11], Flanigan had shown that the variety of complex $3$-dimensional nilpotent associative algebras has an irreducible component which does not contain any rigid algebra — it is instead defined by the closure of a union of a one-parameter family of algebras. Here, we encounter a different situation. Informally, although Theorems G1,  G2 and G3 show that there are generic algebras and generic parametric families in the variety of $3$-dimensional noncommutative Jordan, Kokoris, and standard algebras.
Theorem G1.
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$. It is defined by $5$ rigid algebras and three one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 23. The variety of complex $4$-dimensional nilpotent noncommutative Jordan algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [16, Theorem B].
Theorem G2.
The variety of complex $3$-dimensional Kokoris algebras has dimension $9$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 20. The variety of complex $4$-dimensional nilpotent Kokoris algebras has dimension $13$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in Theorem 21.
Theorem G3.
The variety of complex $3$-dimensional standard algebras has dimension $9$. It is defined by $13$ rigid algebras and one one-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{3}(\mathbb{C})$-orbits of the algebras given in Theorem 22. The variety of complex $4$-dimensional nilpotent standard algebras has dimension $14$. It is defined by $3$ rigid algebras and two one-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in [16, Theorem B].
1. The algebraic classification of non-associative algebras
In our paper, we are working with finite-dimensional vector spaces over the complex field.
1.1. The classification method of noncommutative Jordan algebras
Let $({\bf A},\cdot)$ be an algebra. We consider the following two new products on the underlying vector space ${\bf A}$ defined by
$x\circ y:=\frac{1}{2}(x\cdot y+y\cdot x),\hskip 56.9055pt[x,y]:=\frac{1}{2}(x\cdot y-y\cdot x).$
Let us denote $\textbf{A}^{+}:=(\textbf{A},\circ),$  $\textbf{A}^{-}:=(\textbf{A},[\cdot,\cdot])$ and the associator as $(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z).$
Recall that an algebra $({\bf A},\cdot)$ is called a noncommutative Jordan algebra if it satisfies the identities:
$(x,y,x)=0,$ (1)
$(x\cdot x,y,x)=0.$ (2)
Identity (1) is called the flexibility identity. A flexible algebra ${\bf A}$ satisfies the Jordan identity (2) if and only if the corresponding algebra $\textbf{A}^{+}$ is a Jordan algebra (i.e., a commutative algebra with identity (2)). The class of noncommutative Jordan algebras is extremely large. It contains all Jordan and alternative algebras, as well as all anticommutative algebras.
Definition 1.
An algebra $({\bf P},\circ,[\cdot,\cdot])$ is called a generic Poisson–Jordan algebra (resp., generic Poisson algebra) [^6] To the best of our knowledge, such “associative-commutative” algebras were introduced independently by Cannas and Weinstein under the name “almost Poisson algebras” and by Shestakov under the name “general Poisson algebras” (later changed into “generic Poisson algebras” in a paper by Kolesnikov, Makar-Limanov, and Shestakov). We will use the last terminology. if $({\bf P},\circ)$ is a Jordan (resp., associative commutative) algebra, $({\bf P},[\cdot,\cdot])$ is an anticommutative algebra and these two operations are required to satisfy the following identity:
$[x\circ y,z]=[x,z]\circ y+x\circ[y,z].$
The class of generic Poisson–Jordan algebras is extremely extensive. It contains all Poisson algebras, Malcev–Poisson algebras, and Malcev–Poisson–Jordan algebras, as well as all generic Poisson algebras.
Proposition 2.
$({\bf A},\cdot)$ is a noncommutative Jordan algebra if and only if $({\bf A},\circ,[\cdot,\cdot])$ is a generic Poisson–Jordan algebra.
Proof.
Since the flexible law is equivalent to $[x\circ y,z]=x\circ[y,z]+y\circ[x,z]$, the proof is finished. ∎
Definition 3.
Let $({\bf P}_{1},\circ_{1},[\cdot,\cdot]_{1})$ and $({\bf P}_{2},\circ_{2},[\cdot,\cdot]_{2})$ be two algebras. A linear map $\phi:{\bf P}_{1}\to{\bf P}_{2}$ is a homomorphism if it is preserving the products, that is,
$\phi(x\circ_{1}y)=\phi(x)\circ_{2}\phi(y),\hskip 56.9055pt\phi([x,y]_{1})=[\phi(x),\phi(y)]_{2}.$
Let $({\bf A}_{1},\cdot_{1})$ and $({\bf A}_{2},\cdot_{2})$ be two noncommutative Jordan algebras and let $({\bf A}_{1},\circ_{1},[\cdot,\cdot]_{1})$ and $({\bf A}_{2},\circ_{2},[\cdot,\cdot]_{2})$ be its associated generic Poisson–Jordan algebras. If $({\bf A}_{1},\cdot_{1})$ and $({\bf A}_{2},\cdot_{2})$ are isomorphic, then it is easy to see that the generic Poisson–Jordan algebras $({\bf A}_{1},\circ_{1},[\cdot,\cdot]_{1})$ and $({\bf A}_{2},\circ_{2},[\cdot,\cdot]_{2})$ are isomorphic. Conversely, we can show that if the generic Poisson–Jordan algebras $({\bf A}_{1},\circ_{1},[\cdot,\cdot]_{1})$ and $({\bf A}_{2},\circ_{2},[\cdot,\cdot]_{2})$ are isomorphic, then the noncommutative Jordan algebras $({\bf A}_{1},\cdot_{1})$ and $({\bf A}_{2},\cdot_{2})$ are isomorphic. So we have the following result:
Proposition 4.
Every generic Poisson–Jordan algebra (resp., generic Poisson algebra) $({\bf A},\circ,[\cdot,\cdot])$ is associated with precisely one noncommutative Jordan (resp., Kokoris [^7] About Kokoris algebras, see subsection 1.3.2) algebra $({\bf A},\cdot)$. That is, we have a bijective correspondence between generic Poisson–Jordan (resp., generic Poisson) algebras and noncommutative Jordan (resp., Kokoris) algebras.
Definition 5.
Let $({\bf A},\circ)$ be a Jordan algebra. Let ${\bf Z}^{2}({\bf A},{\bf A})$ be the set of all skew symmetric bilinear maps $\theta:{\bf A}\times{\bf A}\to{\bf A}$ such that
$\theta(x\circ y,z)=\theta(x,z)\circ y+x\circ\theta(y,z).$
For $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$ we define on ${\bf A}$ a product $[\cdot,\cdot]_{\theta}:{\bf A}\times{\bf A}\to{\bf A}$ by
$[x,y]_{\theta}:=\theta(x,y).$ (3)
Lemma 6.
Let $({\bf A},\circ)$ be a Jordan algebra and $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$. Then $({\bf A},\circ,[\cdot,\cdot]_{\theta})$ is a generic Poisson–Jordan algebra endowed with the product defined in (3) and $({\bf A},\cdot_{\theta})$ is a noncommutative Jordan algebra, where $x\cdot_{\theta}y:=x\circ y+[x,y]_{\theta}.$
In the reverse direction, if $({\bf A},\circ,[\cdot,\cdot])$ is a generic Poisson–Jordan algebra, then there exists $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$ such that $({\bf A},\circ,[\cdot,\cdot]_{\theta})$ and $({\bf A},\circ,[\cdot,\cdot])$ are isomorphic. Indeed, consider the skew symmetric bilinear map $\theta:{\bf A}\times{\bf A}\to{\bf A}$ defined by $\theta(x,y):=[x,y].$ Then $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$ and $({\bf A},\circ,[\cdot,\cdot]_{\theta})=({\bf A},\circ,[\cdot,\cdot])$.
Now, let $({\bf A},\circ)$ be a Jordan algebra and $\mathrm{{Aut}({\bf A})}$ be the automorphism group of ${\bf A}$ with respect to product $\circ$. Then $\mathrm{{Aut}({\bf A})}$ acts on ${\bf Z}^{2}({\bf A},{\bf A})$ by
$(\theta\ast\phi)(x,y):=\phi^{-1}\bigl{(}\theta\bigl{(}\phi(x),\phi(y)\bigr{)}\bigr{)},$
where $\phi\in\mathrm{{Aut}({\bf A})}$ and $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$.
Lemma 7.
Let $({\bf A},\circ)$ be a Jordan algebra and $\theta,\vartheta\in{\bf Z}^{2}({\bf A},{\bf A})$. Then the generic Poisson–Jordan algebras $({\bf A},\circ,[\cdot,\cdot]_{\theta})$ and $({\bf A},\circ,[\cdot,\cdot]_{\vartheta})$ are isomorphic if and only if there exists $\phi\in\mathrm{{Aut}({\bf A})}$ satisfying $\theta\ast\phi=\vartheta$.
Proof.
Suppose $\phi:({\bf A},\circ,[\cdot,\cdot]_{\vartheta})\to({\bf A},\circ,[\cdot,\cdot]_{\theta})$ is an isomorphism of generic Poisson–Jordan algebras. Then $\phi\in\text{Aut}({\bf A})$, and by Definition 3, we have $\phi([x,y]_{\vartheta})=[\phi(x),\phi(y)]_{\theta}$. That is, $\phi(\vartheta(x,y))=\theta(\phi(x),\phi(y))$. Hence $\vartheta(x,y)=\phi^{-1}\bigl{(}\theta\bigl{(}\phi(x),\phi(y)\bigr{)}\bigr{)}=(\theta\ast\phi)(x,y)$.
To prove the converse, if $\theta\ast\phi=\vartheta$ then $\phi:({\bf A},\circ,[\cdot,\cdot]_{\vartheta})\to({\bf A},\circ,[\cdot,\cdot]_{\theta})$ is an isomorphism since $\phi(\vartheta(x,y))=\theta(\phi(x),\phi(y))$. ∎
Hence, we have a procedure to classify the generic Poisson–Jordan algebras (and therefore noncommutative Jordan algebras) associated with a given Jordan algebra $({\bf A},\circ)$. It consists of three steps:
  1. Step $1$
    Compute ${\bf Z}^{2}({\bf A},{\bf A})$.
  2. Step $2$
    Find the orbits of $\mathrm{{Aut}({\bf A})}$ on ${\bf Z}^{2}({\bf A},{\bf A})$.
  3. Step $3$
    Choose a representative $\theta$ from each orbit and then construct the generic Poisson–Jordan algebra $({\bf A},\circ,[\cdot,\cdot]_{\theta})$ (the noncommutative Jordan algebra $({\bf A},\cdot_{\theta})$).
Let us introduce the following notations. Let $\{e_{1},\dots,e_{n}\}$ be a fixed basis of a Jordan algebra $({\bf A},\circ)$. Define $\mathrm{\Lambda}^{2}({\bf A},\mathbb{C})$ to be the space of all skew symmetric bilinear forms on ${\bf A}$, that is,
$\mathrm{\Lambda}^{2}({\bf A},\mathbb{C}):=\langle\Delta_{i,j}|1\leq i<j\leq n\rangle$,
where $\Delta_{i,j}$ is the skew-symmetric bilinear form $\Delta_{i,j}:{\bf A}\times{\bf A}\to\mathbb{C}$ defined by
[…]
Now, if $\theta\in{\bf Z}^{2}({\bf A},{\bf A})$ then $\theta$ can be uniquely written as $\theta(x,y)=\sum_{i=1}^{n}B_{i}(x,y)e_{i}$, where $B_{1},\dots,B_{n}$ are skew symmetric bilinear forms on ${\bf A}$. Also, we may write $\theta=(B_{1},\dots,B_{n})$. Let $\phi^{-1}\in\mathrm{Aut}({\bf A})$ be given by the matrix $(b_{ij})$. If $(\theta\ast\phi)(x,y)=\sum_{i=1}^{n}B_{i}^{\prime}(x,y)e_{i}$, then $B_{i}^{\prime}=\sum_{j=1}^{n}b_{ij}\phi^{t}B_{j}\phi$, whenever $i\in\{1,\dots,n\}$.
Remark 8.
Let $X=\begin{pmatrix}\alpha&\beta\end{pmatrix}\in{\mathbb{M}}_{1\times 2}({\mathbb{C}})$ and $X\neq 0$. Then there exists an invertible matrix $A\in{\mathbb{M}}_{2\times 2}({\mathbb{C}})$ such that $XA=\begin{pmatrix}1&0\end{pmatrix}$. To see this, suppose first that $\alpha\neq 0$. Then $\begin{pmatrix}\alpha&\beta\end{pmatrix}\begin{pmatrix}\alpha^{-1}&-\beta\\ 0&\alpha\end{pmatrix}=\begin{pmatrix}1&0\end{pmatrix}$. Assume now that $\alpha=0$. Then $\begin{pmatrix}0&\beta\end{pmatrix}\begin{pmatrix}0&1\\ \beta^{-1}&0\end{pmatrix}=\allowbreak\begin{pmatrix}1&0\end{pmatrix}$.
1.2. The algebraic classification of noncommutative Jordan algebras
Let us mention that by $\circ$ we will denote a commutative multiplication and by [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] we denote an anticommutative multiplication. All products that can be recuperated by commutativity or anticommutativity will be omitted.
1.2.1. The algebraic classification of complex $3$-dimensional noncommutative Jordan algebras
Theorem 9.
(see, [12]) Let ${\bf A}$ be a complex $3$-dimensional associative commutative algebra. Then ${\bf A}$ is isomorphic to one of the following algebras:
${\bf A}_{01}$ :::: $e_{1}\circ e_{1}=e_{2}$
${\bf A}_{02}^{0}$ :::: $e_{1}\circ e_{2}=e_{3}$
${\bf A}_{03}$ :::: $e_{1}\circ e_{1}=e_{2}$ $e_{1}\circ e_{2}=e_{3}$
${\bf A}_{04}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$ $e_{3}\circ e_{3}=e_{3}$
${\bf A}_{05}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$ $e_{2}\circ e_{3}=e_{3}$
${\bf A}_{06}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=e_{2}$ $e_{1}\circ e_{3}=e_{3}$
${\bf A}_{07}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=e_{2}$ $e_{1}\circ e_{3}=e_{3}$ $e_{2}\circ e_{2}=e_{3}$
${\bf A}_{08}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$
${\bf A}_{09}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=e_{2}$
${\bf A}_{10}$ :::: $e_{1}\circ e_{1}=e_{1}$
${\bf A}_{11}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{3}$
Theorem 10.
(see, [12]) Let ${\bf A}$ be a complex $3$-dimensional Jordan algebra. Then ${\bf A}$ is an associative commutative algebra listed in Theorem 9 or isomorphic to one of the following algebras:
${\bf A}_{12}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$ $e_{3}\circ e_{3}=e_{1}+e_{2}$ $e_{1}\circ e_{3}=\frac{1}{2}e_{3}$ $e_{2}\circ e_{3}=\frac{1}{2}e_{3}$
${\bf A}_{13}^{0}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=\frac{1}{2}e_{2}$ $e_{1}\circ e_{3}=e_{3}$
${\bf A}_{14}^{0,0}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=\frac{1}{2}e_{2}$ $e_{1}\circ e_{3}=\frac{1}{2}e_{3}$
${\bf A}_{15}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{3}$ $e_{1}\circ e_{2}=\frac{1}{2}e_{2}$
${\bf A}_{16}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{3}$ $e_{1}\circ e_{2}=\frac{1}{2}e_{2}$ $e_{1}\circ e_{3}=e_{3}$
${\bf A}_{17}^{0}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$ $e_{1}\circ e_{3}=\frac{1}{2}e_{3}$ $e_{2}\circ e_{3}=\frac{1}{2}e_{3}$
${\bf A}_{18}^{0}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{1}\circ e_{2}=\frac{1}{2}e_{2}$
${\bf A}_{19}^{0}$ :::: $e_{1}\circ e_{1}=e_{1}$ $e_{2}\circ e_{2}=e_{2}$ $e_{1}\circ e_{3}=\frac{1}{2}e_{3}$
Theorem 11.
(see, [15]) Let ${\bf A}$ be a complex $3$-dimensional anticommutative algebra. Then ${\bf A}$ is isomorphic to one of the following algebras:
${\bf A}_{20}$ :::: $[e_{2},e_{3}]=e_{1}$
${\bf A}_{21}$ :::: $[e_{1},e_{3}]=e_{1}$ $[e_{2},e_{3}]=e_{2}$
${\bf A}_{22}^{\alpha}$ :::: $[e_{1},e_{3}]=e_{1}+e_{2}$ $[e_{2},e_{3}]=\alpha e_{2}$
${\bf A}_{23}$ :::: $[e_{1},e_{2}]=e_{3}$ $[e_{1},e_{3}]=-e_{2}$ $[e_{2},e_{3}]=e_{1}$
${\bf A}_{24}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3}$ $[e_{1},e_{3}]=e_{1}+e_{3}$ $[e_{2},e_{3}]=\alpha e_{2}$
${\bf A}_{25}$ :::: $[e_{1},e_{2}]=e_{1}$ $[e_{2},e_{3}]=e_{2}$
${\bf A}_{26}$ :::: $[e_{1},e_{2}]=e_{3}$ $[e_{1},e_{3}]=e_{1}$ $[e_{2},e_{3}]=e_{2}$
All listed algebras are non-isomorphic except: ${\bf A}_{22}^{\alpha}\cong{\bf A}_{22}^{\alpha^{-1}},$ ${\bf A}_{24}^{\alpha}\cong{\bf A}_{24}^{\alpha^{-1}}$.
Theorem 12.
Let ${\bf A}$ be a complex $3$-dimensional (noncommutative and non-anticommutative) noncommutative Jordan algebra. Then $\bf{A}$ is isomorphic to one of the following algebras:
${\bf A}_{02}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{2}=\left(1+\alpha\right)e_{3}$ $e_{2}\cdot e_{1}=\left(1-\alpha\right)e_{3}$
${\bf A}_{13}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=(\frac{1}{2}+\alpha)e_{2}$ $e_{2}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{2}$
$e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{14}^{(\alpha,\beta)\neq(0,0)}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=(\frac{1}{2}+\alpha)e_{2}$ $e_{2}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{2}$
$e_{1}\cdot e_{3}=(\frac{1}{2}+\beta)e_{3}$ $e_{3}\cdot e_{1}=(\frac{1}{2}-\beta)e_{3}$
${\bf A}_{17}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{2}=e_{2}$ $e_{1}\cdot e_{3}=(\frac{1}{2}+\alpha)e_{3}$
$e_{3}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{3}$ $e_{2}\cdot e_{3}=(\frac{1}{2}-\alpha)e_{3}$ $e_{3}\cdot e_{2}=(\frac{1}{2}+\alpha)e_{3}$
${\bf A}_{18}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=(\frac{1}{2}+\alpha)e_{2}$ $e_{2}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{2}$
${\bf A}_{19}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{2}=e_{2}$ $e_{1}\cdot e_{3}=(\frac{1}{2}+\alpha)e_{3}$ $e_{3}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{3}$
${\bf A}_{27}$ :::: $e_{1}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{1}=-e_{3}$
${\bf A}_{28}$ :::: $e_{1}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{2}$ $e_{3}\cdot e_{1}=-e_{2}$
${\bf A}_{29}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{2}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$
$e_{3}\cdot e_{1}=e_{3}$ $e_{2}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{2}=-e_{3}$
${\bf A}_{30}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{2}=-e_{3}$
${\bf A}_{31}^{\alpha}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=(\frac{1}{2}+\alpha)e_{2}$ $e_{2}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{2}$
$e_{1}\cdot e_{3}=e_{2}+(\frac{1}{2}+\alpha)e_{3}$ $e_{3}\cdot e_{1}=-e_{2}+(\frac{1}{2}-\alpha)e_{3}$
${\bf A}_{32}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}+e_{3}$ $e_{2}\cdot e_{1}=\frac{1}{2}e_{2}-e_{3}$ $e_{1}\cdot e_{3}=\frac{1}{2}e_{3}$
$e_{3}\cdot e_{1}=\frac{1}{2}e_{3}$ $e_{2}\cdot e_{3}=e_{2}$ $e_{3}\cdot e_{2}=-e_{2}$
${\bf A}_{33}^{\alpha}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$ $e_{2}\cdot e_{1}=\frac{1}{2}e_{2}$
$e_{1}\cdot e_{3}=(\frac{1}{2}+\alpha)e_{3}$ $e_{3}\cdot e_{1}=(\frac{1}{2}-\alpha)e_{3}$
$e_{2}\cdot e_{3}=e_{2}$ $e_{3}\cdot e_{2}=-e_{2}$
All listed algebras are non-isomorphic except: ${\bf A}_{02}^{\alpha}\cong{\bf A}_{02}^{-\alpha}$, ${\bf A}_{17}^{\alpha}\cong{\bf A}_{17}^{-\alpha}$ and ${\bf A}_{14}^{\alpha,\beta}\cong{\bf A}_{14}^{\beta,\alpha}$.
Proof.
By Theorem 10, we may assume ${\bf A}^{+}\in\left\{{\bf A}_{01},{\bf A}_{02}^{0},\ldots,{\bf A}_{19}^{0}\right\}$. So, we are interested in studying $\theta\in{\bf Z}^{2}({\bf A}^{+},{\bf A}^{+}),$ such that $\theta\neq 0.$ We have the following cases.
  1. (1)
    ${\bf A}^{+}\in\left\{{\bf A}_{03},{\bf A}_{04},{\bf A}_{05},{\bf A}_{07},{\bf A}_{08},{\bf A}_{09},{\bf A}_{11},{\bf A}_{12},{\bf A}_{15},{\bf A}_{16}\right\}$. Then ${\bf Z}^{2}\left({\bf A}^{+},{\bf A}^{+}\right)=\{0\}$. In this case, there are no new algebras for us.
  2. (2)
    ${\bf A}^{+}={\bf A}_{01}$. Let $\theta\in{\bf Z}^{2}({\bf A}_{01},{\bf A}_{01})$. Then $\theta=(0,\alpha\Delta_{1,3},\beta\Delta_{1,3})$ for some $\alpha,\beta\in\mathbb{C}$. $\mathrm{Aut}({\bf A}_{01})$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}a_{11}&0&0\\ a_{21}&a_{11}^{2}&a_{23}\\ a_{31}&0&a_{33}\end{pmatrix}.$
    Let $\phi\in\mathrm{Aut}({\bf A}_{01})$. Then $\theta*\phi=(0,\alpha^{\prime}\Delta_{1,3},\beta^{\prime}\Delta_{1,3})$ where
    $\alpha^{\prime}=a_{11}^{-1}(\alpha a_{33}-\beta a_{23}),$  $\beta^{\prime}=\beta a_{11}.$
    Let us distinguish two cases:
    • If $\beta\neq 0$, we define $\phi$ to be the following automorphism:
      $\phi=\begin{pmatrix}{\beta^{-1}}&0&0\\ 0&{\beta^{-2}}&{\alpha}{\beta^{-1}}\\ 0&0&1\end{pmatrix}.$
      Then $\theta\ast\phi=(0,0,\Delta_{1,3})$. So we get the algebra ${\bf A}_{27}$.
    • If $\beta=0$, we define $\phi$ to be the following automorphism:
      $\phi=\begin{pmatrix}\alpha&0&0\\ 0&\alpha^{2}&0\\ 0&0&1\end{pmatrix}.$
      Then $\theta\ast\phi=(0,\Delta_{1,3},0)$. So, we get the algebra ${\bf A}_{28}$.
  3. (3)
    ${\bf A}^{+}={\bf A}_{02}^{0}$. Let $\theta\in{\bf Z}^{2}({\bf A}_{02}^{0},{\bf A}_{02}^{0})$. Then $\theta=(0,0,\alpha\Delta_{1,2})$ for some $\alpha\in\mathbb{C}$. $\mathrm{Aut}({\bf A}_{02}^{0})$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}a_{11}&a_{12}&0\\ a_{21}&a_{22}&0\\ a_{31}&a_{32}&a_{11}a_{22}+a_{12}a_{21}\end{pmatrix},\text{ where }a_{12}=a_{21}=0\text{ or }a_{11}=a_{22}=0.$
    Then $\theta*\phi=(0,0,\alpha^{\prime}\Delta_{1,2})$ and $\alpha^{\prime}=\frac{\alpha a_{11}a_{22}-\alpha a_{12}a_{21}}{a_{11}a_{22}+a_{12}a_{21}}.$ Since $a_{12}=a_{21}=0\text{ or }a_{11}=a_{22}=0$, we have $(\alpha^{\prime})^{2}=\alpha^{2}$. So we get the representatives $\theta^{\alpha}=\left(0,0,\alpha\Delta_{1,2}\right)$. Moreover, we have $\theta^{\alpha}$ and $\theta^{\alpha^{\prime}}$ in the same orbit if and only if $(\alpha^{\prime})^{2}=\alpha^{2}$. So, we get a family of algebras ${\bf A}_{02}^{\alpha\neq 0}$.
  4. (4)
    ${\bf A}^{+}={\bf A}_{06}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{06},{\bf A}_{06}\right)$. Then $\theta=\left(0,\alpha\triangle_{2,3},\beta\triangle_{2,3}\right)$ for some $\alpha,\beta\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{06}\right)$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}1&0&0\\ 0&a_{22}&a_{23}\\ 0&a_{32}&a_{33}\end{pmatrix}.$
    Then $\theta\ast\phi=\left(0,(\alpha a_{33}-\beta a_{23})\triangle_{2,3},(\beta a_{22}-\alpha a_{32})\triangle_{2,3}\right)$ and for suitable $a_{ij},$ we get the representative $\theta=\left(0,0,\triangle_{2,3}\right)$ and the algebra ${\bf A}_{29}$.
  5. (5)
    ${\bf A}^{+}={\bf A}_{10}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{10},{\bf A}_{10}\right)$. Then $\theta=\left(0,\alpha\triangle_{2,3},\beta\triangle_{2,3}\right)$ for some $\alpha,\beta\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{10}\right)$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}1&0&0\\ 0&a_{22}&a_{23}\\ 0&a_{32}&a_{33}\end{pmatrix}.$
    Then $\theta\ast\phi=\left(0,(\alpha a_{33}-\beta a_{23})\triangle_{2,3},(\beta a_{22}-\alpha a_{32})\triangle_{2,3}\right)$ and for suitable $a_{ij},$ we get the representative $\theta=\left(0,0,\triangle_{2,3}\right)$ and the algebra ${\bf A}_{30}$.
  6. (6)
    ${\bf A}^{+}={\bf A}_{13}^{0}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{13}^{0},{\bf A}_{13}^{0}\right)$. Then $\theta=\left(0,\alpha\Delta_{1,2},0\right)$ for some $\alpha\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{13}^{0}\right)$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}1&0&0\\ a_{21}&a_{22}&0\\ 0&0&a_{33}\end{pmatrix}.$
    Then $\theta\ast\phi=\left(0,\alpha^{\prime}\Delta_{1,2},0\right)$ and $\alpha^{\prime}=\alpha$. Thus we have the the family of representatives $\theta^{\alpha}=\left(0,\alpha\Delta_{1,2},0\right)$ and the family of algebras ${\bf A}_{13}^{\alpha\neq 0}$.
  7. (7)
    ${\bf A}^{+}={\bf A}_{14}^{0,0}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{14}^{0,0},{\bf A}_{14}^{0,0}\right)$. Then
    $\theta=\left(0,\alpha_{1}\Delta_{1,2}+\alpha_{2}\Delta_{1,3}+\alpha_{3}\Delta_{2,3},\alpha_{4}\Delta_{1,2}+\alpha_{5}\Delta_{1,3}+\alpha_{6}\Delta_{2,3}\right)$
    for some $\alpha_{1},\ldots,\alpha_{6}\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{14}^{0,0}\right)$ consists of the invertible matrices of the following form:
    $\phi=\begin{pmatrix}1&0&0\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{pmatrix}.$
    Then $\theta\ast\phi=\left(0,\beta_{1}\Delta_{1,2}+\beta_{2}\Delta_{1,3}+\beta_{3}\Delta_{2,3},\beta_{4}\Delta_{1,2}+\beta_{5}\Delta_{1,3}+\beta_{6}\Delta_{2,3}\right)$ and
    $\beta_{1}$ $=$ $\frac{(\alpha_{1}a_{22}+\alpha_{2}a_{32}+\alpha_{3}(a_{21}a_{32}-a_{22}a_{31}))a_{33}-(\alpha_{4}a_{22}+\alpha_{5}a_{32}+\alpha_{6}(a_{21}a_{32}-a_{22}a_{31}))a_{23}}{\det\phi},$
    $\beta_{2}$ $=$ $\frac{(\alpha_{1}a_{23}+\alpha_{2}a_{33}+\alpha_{3}(a_{21}a_{33}-a_{31}a_{23}))a_{33}-(\alpha_{4}a_{23}+\alpha_{5}a_{33}+\alpha_{6}(a_{21}a_{33}-a_{31}a_{23}))a_{23}}{\det\phi},$
    $\beta_{3}$ $=$ $\alpha_{3}a_{33}-\alpha_{6}a_{23},$
    $\beta_{4}$ $=$ $\frac{-(\alpha_{1}a_{22}+\alpha_{2}a_{32}+\alpha_{3}(a_{21}a_{32}-a_{22}a_{31}))a_{32}+(\alpha_{4}a_{22}+\alpha_{5}a_{32}+\alpha_{6}(a_{21}a_{32}-a_{22}a_{31}))a_{22}}{\det\phi},$
    $\beta_{5}$ $=$ $\frac{-(\alpha_{1}a_{23}+\alpha_{2}a_{33}+\alpha_{3}(a_{21}a_{33}-a_{31}a_{23}))a_{32}+(\alpha_{4}a_{23}+\alpha_{5}a_{33}+\alpha_{6}(a_{21}a_{33}-a_{31}a_{23}))a_{22}}{\det\phi},$
    $\beta_{6}$ $=$ $\alpha_{6}a_{22}-\alpha_{3}a_{32}.$
    By Remark 8, we may assume $\left(\alpha_{3},\alpha_{6}\right)\in\left\{\left(0,0\right),\left(1,0\right)\right\}$. Let us consider these cases separately.
    • Assume first that $(\alpha_{3},\alpha_{6})=(0,0)$. Then
      $\begin{pmatrix}\beta_{1}&\beta_{2}\\ \beta_{4}&\beta_{5}\end{pmatrix}=\left(\begin{array}[]{cc}a_{22}&a_{23}\\ a_{32}&a_{33}\end{array}\right)^{-1}\begin{pmatrix}\alpha_{1}&\alpha_{2}\\ \alpha_{4}&\alpha_{5}\end{pmatrix}\left(\begin{array}[]{cc}a_{22}&a_{23}\\ a_{32}&a_{33}\end{array}\right).$
      From here, we may assume $\begin{pmatrix}\alpha_{1}&\alpha_{2}\\ \alpha_{4}&\alpha_{5}\end{pmatrix}\in\left\{\begin{pmatrix}\alpha&0\\ 0&\beta\end{pmatrix},\begin{pmatrix}\alpha&1\\ 0&\alpha\end{pmatrix}\right\}$. So we get representatives
      $\theta_{1}^{\alpha,\beta}=\left(0,\alpha\Delta_{1,2},\beta\Delta_{1,3}\right)$ and $\theta_{2}^{\alpha}=\left(0,\alpha\Delta_{1,2}+\Delta_{1,3},\alpha\Delta_{1,3}\right)$.
      Clearly, $\theta_{1}^{\alpha,\beta}$ and $\theta_{2}^{\alpha}$ are not in the same orbit. Furthermore, we have $\theta_{1}^{\alpha,\beta}$ and $\theta_{1}^{\alpha^{\prime},\beta^{\prime}}$ in the same orbit if and only if $\left\{\alpha,\beta\right\}=\left\{\alpha^{\prime},\beta^{\prime}\right\}$. Also, $\theta_{2}^{\alpha}$ and $\theta_{2}^{\alpha^{\prime}}$ are in the same orbit if and only if $\alpha=\alpha^{\prime}$. Thus we get families of algebras ${\bf A}_{14}^{(\alpha,\beta)\neq(0,0)}$ and ${\bf A}_{31}^{\alpha}$.
    • Assume now that $(\alpha_{3},\alpha_{6})=(1,0)$. Then we have the following cases:
      • $\alpha_{4}\neq 0$. We define $\phi$ to be the following automorphism:
        $\phi=\begin{pmatrix}1&0&0\\ -\frac{\alpha_{5}^{2}+\alpha_{2}\alpha_{4}}{\alpha_{4}}&\frac{1}{\alpha_{4}}&-\frac{\alpha_{5}}{\alpha_{4}}\\ \alpha_{1}+\alpha_{5}&0&1\end{pmatrix}.$
        Then $\theta\ast\phi=\left(0,\Delta_{23},\Delta_{12}\right)$. So we get the algebra ${\bf A}_{32}$.
      • $\alpha_{4}=0$. We choose $\phi$ to be the following automorphism:
        $\phi=\begin{pmatrix}1&0&0\\ -\alpha_{2}&1&0\\ \alpha_{1}&0&1\end{pmatrix}.$
        Then $\theta\ast\phi=\left(0,\Delta_{23},\alpha_{5}\Delta_{13}\right)$. Hence we get the representatives $\theta^{\alpha}=\left(0,\Delta_{23},\alpha\Delta_{13}\right)$. The representatives $\theta^{\alpha}$ and $\theta^{\beta}$ are in the same orbit if and only if $\alpha=\beta$. Thus, we obtain the algebras ${\bf A}_{33}^{\alpha}$.
  8. (8)
    ${\bf A}^{+}={\bf A}_{17}^{0}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{17}^{0},{\bf A}_{17}^{0}\right)$. Then $\theta=\left(0,0,\alpha\left(\Delta_{1,3}-\Delta_{2,3}\right)\right)$ for some $\alpha\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{17}^{0}\right)$ consists of the invertible matrices of the following form:
    $\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ a_{31}&-a_{31}&a_{33}\end{array}\right)\mbox{ and }\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ a_{31}&-a_{31}&a_{33}\end{array}\right).$
    Then $\theta\ast\phi=\left(0,0,\alpha^{\prime}\left(\Delta_{1,3}-\Delta_{2,3}\right)\right)$ and $(\alpha^{\prime})^{2}=\alpha^{2}$. So we get the family of algebras ${\bf A}_{17}^{\alpha\neq 0}$.
  9. (9)
    ${\bf A}^{+}={\bf A}_{18}^{0}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{18}^{0},{\bf A}_{18}^{0}\right)$. Then $\theta=\left(0,\alpha\Delta_{1,2},0\right)$ for some $\alpha\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{18}^{0}\right)$ consists of the invertible matrices of the following form
    $\phi=\left(\begin{array}[]{ccc}1&0&0\\ a_{21}&a_{22}&0\\ 0&0&a_{33}\end{array}\right).$
    Then $\theta\ast\phi=\left(0,\alpha^{\prime}\Delta_{1,2},0\right)$ and $\alpha^{\prime}=\alpha$. So we get the family of algebras ${\bf A}_{18}^{\alpha\neq 0}$.
  10. (10)
    ${\bf A}^{+}={\bf A}_{19}^{0}$. Let $\theta\in{\bf Z}^{2}\left({\bf A}_{19}^{0},{\bf A}_{19}^{0}\right)$. Then $\theta=\left(0,0,\alpha\Delta_{1,3}\right)$ for some $\alpha\in\mathbb{C}$. $\mathrm{Aut}\left({\bf A}_{19}^{0}\right)$ consists of the invertible matrices of the following form:
    $\phi=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ a_{31}&0&a_{33}\end{array}\right).$
    Then $\theta\ast\phi=\left(0,0,\alpha^{\prime}\Delta_{1,3}\right)$ and $\alpha^{\prime}=\alpha$. So we get the family of algebras ${\bf A}_{19}^{\alpha\neq 0}$.
1.2.2. The algebraic classification of complex $4$-dimensional nilpotent noncommutative Jordan algebras
Let us remember the classification of $4$-dimensional nilpotent noncommutative Jordan algebras obtained in [16]. This classification will be useful for results given below, in subsections 1.3.3 and 1.4.2.
Theorem 13.
Let $\mathcal{J}$ be a complex $4$-dimensional nilpotent noncommutative Jordan algebra. Then $\mathcal{J}$ is a $2$-step nilpotent algebra listed in [20] or isomorphic to one of the following algebras:
${\mathcal{J}}_{01}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{3}$ $e_{2}e_{1}=e_{3}$
${\mathcal{J}}_{02}$ :::: $e_{1}e_{1}=e_{2}$ $e_{2}e_{3}=e_{4}$ $e_{3}e_{2}=e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{03}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{1}=e_{4}$ $e_{3}e_{3}=e_{4}$
${\mathcal{J}}_{04}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{05}$ :::: $e_{1}e_{1}=e_{2}$ $e_{2}e_{3}=e_{4}$ $e_{3}e_{2}=e_{4}$
${\mathcal{J}}_{06}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{3}=e_{4}$
${\mathcal{J}}_{07}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}+e_{4}$ $e_{2}e_{3}=e_{4}$ $e_{3}e_{1}=e_{4}$ $e_{3}e_{2}=e_{4}$
${\mathcal{J}}_{08}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}+e_{4}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{09}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}+e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{10}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}$ $e_{2}e_{3}=e_{4}$ $e_{3}e_{1}=e_{4}$ $e_{3}e_{2}=e_{4}$
${\mathcal{J}}_{11}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{12}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{13}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{14}$ :::: $e_{1}e_{1}=e_{4}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{15}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}+e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{16}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{17}$ :::: $e_{1}e_{1}=e_{4}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{18}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=e_{4}$
1.3. The algebraic classification of Kokoris algebras
A flexible algebra $\bf A$ is called a Kokoris algebra if $\bf A^{+}$ is associative. Kokoris algebras were introduced in [24] and they also appear in [40, 38, 7]. $({\bf A},\cdot)$ is a Kokoris algebra if and only if $({\bf A},\circ,[\cdot,\cdot])$ is a generic Poisson algebra. Also, observe that Poisson algebras are related to Lie admissible Kokoris algebras.
1.3.1. The algebraic classification of complex $3$-dimensional Kokoris algebras
Theorem 14.
Let $\bf{K}$ be a complex $3$-dimensional Kokoris algebra. Then $\bf{K}$ is an associative commutative algebra listed in Theorem 9, an anticommutative algebra listed in Theorem 11 or it is isomorphic to one of the following algebras:
${\bf A}_{02}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{2}=\left(1+\alpha\right)e_{3}$ $e_{2}\cdot e_{1}=\left(1-\alpha\right)e_{3}$
${\bf A}_{27}$ :::: $e_{1}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{1}=-e_{3}$
${\bf A}_{28}$ :::: $e_{1}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{2}$ $e_{3}\cdot e_{1}=-e_{2}$
${\bf A}_{29}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{2}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$
$e_{3}\cdot e_{1}=e_{3}$ $e_{2}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{2}=-e_{3}$
${\bf A}_{30}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{2}=-e_{3}$
All listed algebras are non-isomorphic except: ${\bf A}_{02}^{\alpha}\cong{\bf A}_{02}^{-\alpha}.$
1.3.2. The algebraic classification of complex $3$-dimensional flexible power-associative algebras
It is a well known fact that if an algebra A is power-associative then $\textbf{A}^{+}$ is power-associative. For flexible algebras the converse is also true [22]. In [34], it is proved that any commutative power-associative algebra with dimension at most $3$ is a Jordan algebra. Since a flexible algebra A is a noncommutative Jordan algebra if and only if the corresponding algebra $\textbf{A}^{+}$ is a Jordan algebra, all flexible power-associative algebras of dimension at most $3$ are noncommutative Jordan algebras.
Corollary 15.
Let ${\bf A}$ be a complex $3$-dimensional flexible power-associative algebra, then ${\bf A}$ is isomorphic to one algebra listed in Theorems 9, 10, 11 and 12.
1.3.3. The algebraic classification of complex $4$-dimensional nilpotent Kokoris algebras
Theorem 16.
Let $\mathcal{J}$ be a complex $4$-dimensional nilpotent Kokoris algebra. Then $\mathcal{J}$ is a $2$-step nilpotent algebra listed in [20] or isomorphic to one of the following algebras:
${\mathcal{J}}_{01}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{3}$ $e_{2}e_{1}=e_{3}$
${\mathcal{J}}_{03}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{1}=e_{4}$ $e_{3}e_{3}=e_{4}$
${\mathcal{J}}_{04}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{1}=e_{4}$
${\mathcal{J}}_{06}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{4}$ $e_{2}e_{1}=e_{4}$ $e_{3}e_{3}=e_{4}$
${\mathcal{J}}_{13}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{14}$ :::: $e_{1}e_{1}=e_{4}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{15}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}+e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{16}$ :::: $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{17}$ :::: $e_{1}e_{1}=e_{4}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=-e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=-e_{4}$
${\mathcal{J}}_{18}$ :::: $e_{1}e_{1}=e_{2}$ $e_{1}e_{2}=e_{3}$ $e_{1}e_{3}=e_{4}$ $e_{2}e_{1}=e_{3}$ $e_{2}e_{2}=e_{4}$ $e_{3}e_{1}=e_{4}$
1.4. The algebraic classification of standard algebras
The notions of standard algebras were introduced in [2]. An algebra is defined to be standard in case the following two identities hold:
$(x,y,z)+(z,x,y)-(x,z,y)$ $=$ $0,$
$(x,y,wz)+(w,y,xz)+(z,y,wx)$ $=$ $0$
Standard algebras include all associative algebras and Jordan algebras. It is proved that the variety of standard algebras is just the minimal variety containing the variety of associative algebras and the variety of Jordan algebras [6]. Moreover, every standard algebra is a noncommutative Jordan algebra and is therefore power-associative. By some direct verification of standard identities in noncommutative Jordan algebras, we have the following statements.
1.4.1. The algebraic classification of complex $3$-dimensional standard algebras
Theorem 17.
Let $\bf{S}$ be a complex $3$-dimensional standard algebra. Then $\bf{S}$ is a Jordan algebra listed in Theorems 9 and 10 or isomorphic to one of the following algebras:
${\bf A}_{02}^{\alpha\neq 0}$ :::: $e_{1}\cdot e_{2}=\left(1+\alpha\right)e_{3}$ $e_{2}\cdot e_{1}=\left(1-\alpha\right)e_{3}$
${\bf A}_{13}^{\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{13}^{-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{14}^{0,\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$ $e_{2}\cdot e_{1}=\frac{1}{2}e_{2}$ $e_{1}\cdot e_{3}=e_{3}$
${\bf A}_{14}^{0,-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=\frac{1}{2}e_{2}$ $e_{2}\cdot e_{1}=\frac{1}{2}e_{2}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{14}^{\frac{1}{2},\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$
${\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{1}=e_{2}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{3}\cdot e_{1}=e_{3}$
${\bf A}_{17}^{\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{2}=e_{2}$ $e_{1}\cdot e_{3}=e_{3}$ $e_{3}\cdot e_{2}=e_{3}$
${\bf A}_{18}^{\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$
${\bf A}_{18}^{-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{1}=e_{2}$
${\bf A}_{19}^{\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{1}\cdot e_{2}=e_{2}$ $e_{3}\cdot e_{3}=e_{3}$
${\bf A}_{19}^{-\frac{1}{2}}$ :::: $e_{1}\cdot e_{1}=e_{1}$ $e_{2}\cdot e_{1}=e_{2}$ $e_{3}\cdot e_{3}=e_{3}$
${\bf A}_{20}$ :::: $e_{2}\cdot e_{3}=e_{1}$ $e_{3}\cdot e_{2}=-e_{1}$
${\bf A}_{28}$ :::: $e_{1}\cdot e_{1}=e_{2}$ $e_{1}\cdot e_{3}=e_{2}$ $e_{3}\cdot e_{1}=-e_{2}$
All listed algebras are non-isomorphic except: ${\bf A}_{02}^{\alpha}\cong{\bf A}_{02}^{-\alpha}.$
1.4.2. The algebraic classification of complex $4$-dimensional nilpotent standard algebras
Theorem 18.
Any complex $4$-dimensional nilpotent noncommutative Jordan algebra, listed in Theorem 13, is a standard algebra.
2. The geometric classification of non-associative algebras
2.1. Definitions and notation
Given an $n$-dimensional vector space $\mathbb{V}$, the set ${\rm Hom}(\mathbb{V}\otimes\mathbb{V},\mathbb{V})\cong\mathbb{V}^{*}\otimes\mathbb{V}^{*}\otimes\mathbb{V}$ is a vector space of dimension $n^{3}$. This space has the structure of the affine variety $\mathbb{C}^{n^{3}}.$ Indeed, let us fix a basis $e_{1},\dots,e_{n}$ of $\mathbb{V}$. Then any $\mu\in{\rm Hom}(\mathbb{V}\otimes\mathbb{V},\mathbb{V})$ is determined by $n^{3}$ structure constants $c_{ij}^{k}\in\mathbb{C}$ such that $\mu(e_{i}\otimes e_{j})=\sum\limits_{k=1}^{n}c_{ij}^{k}e_{k}$. A subset of ${\rm Hom}(\mathbb{V}\otimes\mathbb{V},\mathbb{V})$ is Zariski-closed if it can be defined by a set of polynomial equations in the variables $c_{ij}^{k}$ ($1\leq i,j,k\leq n$).
Let $T$ be a set of polynomial identities. The set of algebra structures on $\mathbb{V}$ satisfying polynomial identities from $T$ forms a Zariski-closed subset of the variety ${\rm Hom}(\mathbb{V}\otimes\mathbb{V},\mathbb{V})$. We denote this subset by $\mathbb{L}(T)$. The general linear group ${\rm GL}(\mathbb{V})$ acts on $\mathbb{L}(T)$ by conjugations:
$(g*\mu)(x\otimes y)=g\mu(g^{-1}x\otimes g^{-1}y)$
for $x,y\in\mathbb{V}$, $\mu\in\mathbb{L}(T)\subset{\rm Hom}(\mathbb{V}\otimes\mathbb{V},\mathbb{V})$ and $g\in{\rm GL}(\mathbb{V})$. Thus, $\mathbb{L}(T)$ is decomposed into ${\rm GL}(\mathbb{V})$-orbits that correspond to the isomorphism classes of algebras. Let ${\mathcal{O}}(\mu)$ denote the orbit of $\mu\in\mathbb{L}(T)$ under the action of ${\rm GL}(\mathbb{V})$ and $\overline{{\mathcal{O}}(\mu)}$ denote the Zariski closure of ${\mathcal{O}}(\mu)$.
Let $\bf A$ and $\bf B$ be two $n$-dimensional algebras satisfying the identities from $T$, and let $\mu,\lambda\in\mathbb{L}(T)$ represent $\bf A$ and $\bf B$, respectively. We say that $\bf A$ degenerates to $\bf B$ and write $\bf A\to\bf B$ if $\lambda\in\overline{{\mathcal{O}}(\mu)}$. Note that in this case we have $\overline{{\mathcal{O}}(\lambda)}\subset\overline{{\mathcal{O}}(\mu)}$. Hence, the definition of degeneration does not depend on the choice of $\mu$ and $\lambda$. If $\bf A\not\cong\bf B$, then the assertion $\bf A\to\bf B$ is called a proper degeneration. We write $\bf A\not\to\bf B$ if $\lambda\not\in\overline{{\mathcal{O}}(\mu)}$.
Let $\bf A$ be represented by $\mu\in\mathbb{L}(T)$. Then $\bf A$ is rigid in $\mathbb{L}(T)$ if ${\mathcal{O}}(\mu)$ is an open subset of $\mathbb{L}(T)$. Recall that a subset of a variety is called irreducible if it cannot be represented as a union of two non-trivial closed subsets. A maximal irreducible closed subset of a variety is called an irreducible component. It is well known that any affine variety can be represented as a finite union of its irreducible components in a unique way. The algebra $\bf A$ is rigid in $\mathbb{L}(T)$ if and only if $\overline{{\mathcal{O}}(\mu)}$ is an irreducible component of $\mathbb{L}(T)$.
Method of the description of degenerations of algebras. In the present work we use the methods applied to Lie algebras in [13]. First of all, if $\bf A\to\bf B$ and $\bf A\not\cong\bf B$, then $\mathfrak{Der}(\bf A)<\mathfrak{Der}(\bf B)$, where $\mathfrak{Der}(\bf A)$ is the algebra of derivations of $\bf A$. We compute the dimensions of algebras of derivations and check the assertion $\bf A\to\bf B$ only for such $\bf A$ and $\bf B$ that $\mathfrak{Der}(\bf A)<\mathfrak{Der}(\bf B)$.
To prove degenerations, we construct families of matrices parametrized by $t$. Namely, let $\bf A$ and $\bf B$ be two algebras represented by the structures $\mu$ and $\lambda$ from $\mathbb{L}(T)$ respectively. Let $e_{1},\dots,e_{n}$ be a basis of $\mathbb{V}$ and $c_{ij}^{k}$ ($1\leq i,j,k\leq n$) be the structure constants of $\lambda$ in this basis. If there exist $a_{i}^{j}(t)\in\mathbb{C}$ ($1\leq i,j\leq n$, $t\in\mathbb{C}^{*}$) such that $E_{i}^{t}=\sum\limits_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of $\mathbb{V}$ for any $t\in\mathbb{C}^{*}$, and the structure constants of $\mu$ in the basis $E_{1}^{t},\dots,E_{n}^{t}$ are such rational functions $c_{ij}^{k}(t)\in\mathbb{C}[t]$ that $c_{ij}^{k}(0)=c_{ij}^{k}$, then $\bf A\to\bf B$. In this case $E_{1}^{t},\dots,E_{n}^{t}$ is called a parametrized basis for $\bf A\to\bf B$. In case of $E_{1}^{t},E_{2}^{t},\ldots,E_{n}^{t}$ is a parametric basis for ${\bf A}\to{\bf B},$ it will be denoted by ${\bf A}\xrightarrow{(E_{1}^{t},E_{2}^{t},\ldots,E_{n}^{t})}{\bf B}$. To simplify our equations, we will use the notation $A_{i}=\langle e_{i},\dots,e_{n}\rangle,\ i=1,\ldots,n$ and write simply $A_{p}A_{q}\subset A_{r}$ instead of $c_{ij}^{k}=0$ ($i\geq p$, $j\geq q$, $k<r$).
Let ${\bf A}(*):=\{{\bf A}(\alpha)\}_{\alpha\in I}$ be a series of algebras, and let $\bf B$ be another algebra. Suppose that for $\alpha\in I$, $\bf A(\alpha)$ is represented by the structure $\mu(\alpha)\in\mathbb{L}(T)$ and $\bf B$ is represented by the structure $\lambda\in\mathbb{L}(T)$. Then we say that $\bf A(*)\to\bf B$ if $\lambda\in\overline{\{{\mathcal{O}}(\mu(\alpha))\}_{\alpha\in I}}$, and $\bf A(*)\not\to\bf B$ if $\lambda\not\in\overline{\{{\mathcal{O}}(\mu(\alpha))\}_{\alpha\in I}}$.
Let $\bf A(*)$, $\bf B$, $\mu(\alpha)$ ($\alpha\in I$) and $\lambda$ be as above. To prove $\bf A(*)\to\bf B$ it is enough to construct a family of pairs $(f(t),g(t))$ parametrized by $t\in\mathbb{C}^{*}$, where $f(t)\in I$ and $g(t)\in{\rm GL}(\mathbb{V})$. Namely, let $e_{1},\dots,e_{n}$ be a basis of $\mathbb{V}$ and $c_{ij}^{k}$ ($1\leq i,j,k\leq n$) be the structure constants of $\lambda$ in this basis. If we construct $a_{i}^{j}:\mathbb{C}^{*}\to\mathbb{C}$ ($1\leq i,j\leq n$) and $f:\mathbb{C}^{*}\to I$ such that $E_{i}^{t}=\sum\limits_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of $\mathbb{V}$ for any $t\in\mathbb{C}^{*}$, and the structure constants of $\mu({f(t)})$ in the basis $E_{1}^{t},\dots,E_{n}^{t}$ are such rational functions $c_{ij}^{k}(t)\in\mathbb{C}[t]$ that $c_{ij}^{k}(0)=c_{ij}^{k}$, then $\bf A(*)\to\bf B$. In this case $E_{1}^{t},\dots,E_{n}^{t}$ and $f(t)$ are called a parametrized basis and a parametrized index for $\bf A(*)\to\bf B$, respectively.
We now explain how to prove $\bf A(*)\not\to\mathcal{\bf}B$. Note that if $\mathfrak{Der}\ \bf A(\alpha)>\mathfrak{Der}\ \bf B$ for all $\alpha\in I$ then $\bf A(*)\not\to\bf B$. One can also use the following Lemma, whose proof is the same as the proof of Lemma 1.5 from [13].
Lemma 19.
Let $\mathfrak{B}$ be a Borel subgroup of ${\rm GL}(\mathbb{V})$ and $\mathcal{R}\subset\mathbb{L}(T)$ be a $\mathfrak{B}$-stable closed subset. If $\bf A(*)\to\bf B$ and for any $\alpha\in I$ the algebra $\bf A(\alpha)$ can be represented by a structure $\mu(\alpha)\in\mathcal{R}$, then there is $\lambda\in\mathcal{R}$ representing $\bf B$.
2.2. The geometric classification of Kokoris algebras
The main result of the present section is the following theorems.
Theorem 20.
The variety of complex $3$-dimensional Kokoris algebras has dimension $9$ and it has $5$ irreducible components defined by
$\mathcal{C}_{1}=\overline{\mathcal{O}({\bf A}_{02}^{\alpha})},$  $\mathcal{C}_{2}=\overline{\mathcal{O}({\bf A}_{04})},$  $\mathcal{C}_{3}=\overline{\mathcal{O}({\bf A}_{24}^{\alpha})},$  $\mathcal{C}_{4}=\overline{\mathcal{O}({\bf A}_{29})},$  $\mathcal{C}_{5}=\overline{\mathcal{O}({\bf A}_{30})}.$
In particular, there are only three rigid algebras in this variety.
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}({\bf A}_{04})=\dim\mathcal{O}({\bf A}_{24}^{\alpha})$ $=$ $9,$
$\dim\mathcal{O}({\bf A}_{29})=\dim\mathcal{O}({\bf A}_{30})$ $=$ $7,$
$\dim\mathcal{O}({\bf A}_{02}^{\alpha})$ $=$ $6$
${\bf A}_{04}$ is commutative and ${\bf A}_{24}^{\alpha}$ is anticommutative, hence $\{{\bf A}_{04},\ {\bf A}_{24}^{\alpha}\}\not\to\{{\bf A}_{02}^{\alpha},\ {\bf A}_{29},\ {\bf A}_{30}\}.$ The principal non-degenerations $\{{\bf A}_{29},\ \ {\bf A}_{30}\}\not\to{\bf A}_{02}^{\alpha}$ are justified by the following condition
=$\mathcal{R}=\left\{c_{12}^{3}=-c_{21}^{3},\ c_{13}^{2}=c_{31}^{2}=0,\ c_{23}^{1}=c_{32}^{1}=0\right\}.$
All necessary degenerations are given below
  1. (1)
    ${\bf A}_{30}\xrightarrow{(te_{1}+e_{2}-e_{3},-te_{2}+te_{3},t^{2}e_{3})}{\bf A}_{27},$  ${\bf A}_{30}\xrightarrow{(te_{1}+e_{2},-te_{2},-te_{3})}{\bf A}_{28}.$
  2. (2)
    Thanks to [15], ${\bf A}_{24}^{\alpha}\to{\bf A}_{20},\ {\bf A}_{21},\ {\bf A}_{22}^{\alpha},\ {\bf A}_{23},\ {\bf A}_{25},\ {\bf A}_{26}.$
  3. (3)
    Thanks to [12], ${\bf A}_{04}\to{\bf A}_{01},\ {\bf A}_{03},\ {\bf A}_{05},\ {\bf A}_{06},\ {\bf A}_{07},\ {\bf A}_{08},\ {\bf A}_{09},\ {\bf A}_{10},\ {\bf A}_{11}.$
Theorem 21.
The variety of complex $4$-dimensional nilpotent Kokoris algebras has dimension $13$ and it has $5$ irreducible components defined by
$\mathcal{C}_{1}=\overline{\mathcal{O}({\mathcal{J}}_{03})},$  $\mathcal{C}_{2}=\overline{\mathcal{O}({\mathcal{J}}_{17})},$  $\mathcal{C}_{3}=\overline{\mathcal{O}({\mathcal{J}}_{18})},$ $\mathcal{C}_{3}=\overline{\mathcal{O}(\mathfrak{N}_{2}(\alpha))},$  $\mathcal{C}_{5}=\overline{\mathcal{O}(\mathfrak{N}_{3}(\alpha))}.$
In particular, there are only three rigid algebras in this variety.
Proof.
All necessary degenerations are given below
  1. (1)
    ${\mathcal{J}}_{03}\xrightarrow{(t^{-2}e_{1},t^{-4}e_{2},t^{-3}e_{3},t^{-6}e_{4})}{\mathcal{J}}_{06},$ ${\mathcal{J}}_{18}\xrightarrow{(e_{1},e_{2},e_{3},t^{-1}e_{4})}{\mathcal{J}}_{01}.$
  2. (2)
    Thanks to [16], ${\mathcal{J}}_{17}\to{\mathcal{J}}_{14},\ {\mathcal{J}}_{15},\ {\mathcal{J}}_{16};$ ${\mathcal{J}}_{03}\to{\mathcal{J}}_{04};$ ${\mathcal{J}}_{14}\to{\mathcal{J}}_{13}.$
Recall that the geometric classification of $4$-dimensional noncommutative Jordan algebras was given in [16]. Hence, ${\mathcal{J}}_{17},$ ${\mathcal{J}}_{18}$ and
$\begin{array}[]{lllllll}\mathfrak{N}_{2}(\alpha)&e_{1}e_{1}=e_{3},&e_{1}e_{2}=e_{4},&e_{2}e_{1}=-\alpha e_{3},&e_{2}e_{2}=-e_{4}\\ \mathfrak{N}_{3}(\alpha)&e_{1}e_{1}=e_{4},&e_{1}e_{2}=\alpha e_{4},&e_{2}e_{1}=-\alpha e_{4},&e_{2}e_{2}=e_{4},&e_{3}e_{3}=e_{4},\end{array}$
also, give irreducible components in the variety of $4$-dimensional nilpotent Kokoris algebras. Since, $\mathfrak{Der}({\mathcal{J}}_{03})=3$ and $\mathfrak{Der}({\mathcal{J}}_{17})=4$, we have ${\mathcal{J}}_{17}\not\to{\mathcal{J}}_{03}.$ Since, ${\mathcal{J}}_{18}$ is commutative and ${\mathcal{J}}_{03}$ is noncommutative, we have ${\mathcal{J}}_{18}\not\to{\mathcal{J}}_{03}.$
2.3. The geometric classification of standard algebras
The main result of the present section is the following theorem.
Theorem 22.
The variety of complex $3$-dimensional standard algebras has dimension $9$ and it has $14$ irreducible components defined by
$\mathcal{C}_{1}=\overline{\mathcal{O}({\bf A}_{02}^{\alpha})},$  $\mathcal{C}_{2}=\overline{\mathcal{O}({\bf A}_{04})},$  $\mathcal{C}_{3}=\overline{\mathcal{O}({\bf A}_{12})},$  $\mathcal{C}_{4}=\overline{\mathcal{O}({\bf A}_{14}^{0,0})},$  $\mathcal{C}_{5}=\overline{\mathcal{O}({\bf A}_{14}^{\frac{1}{2},\frac{1}{2}})},$
$\mathcal{C}_{6}=\overline{\mathcal{O}({\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}})},$  $\mathcal{C}_{7}=\overline{\mathcal{O}({\bf A}_{14}^{0,\frac{1}{2}})},$  $\mathcal{C}_{8}=\overline{\mathcal{O}({\bf A}_{14}^{0,-\frac{1}{2}})},$  $\mathcal{C}_{9}=\overline{\mathcal{O}({\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}})},$ $\mathcal{C}_{10}=\overline{\mathcal{O}({\bf A}_{16})},$
$\mathcal{C}_{11}=\overline{\mathcal{O}({\bf A}_{17}^{\frac{1}{2}})},$  $\mathcal{C}_{12}=\overline{\mathcal{O}({\bf A}_{19}^{0})},$  $\mathcal{C}_{13}=\overline{\mathcal{O}({\bf A}_{19}^{\frac{1}{2}})},$  $\mathcal{C}_{14}=\overline{\mathcal{O}({\bf A}_{19}^{-\frac{1}{2}})}.$
In particular, there are only thirteen rigid algebras in this variety.
Proof.
Thanks to [12], we have a description of all degenerations between $3$-dimensional Jordan algebras. Hence, algebras ${\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{14}^{0,0},$ ${\bf A}_{16}$ and ${\bf A}_{19}^{0}$ are rigid in the variety of $3$-dimensional Jordan algebras, but algebras ${\bf A}_{01},$ ${\bf A}_{03},$ ${\bf A}_{05},$ ${\bf A}_{06},$ ${\bf A}_{07},$ ${\bf A}_{08},$ ${\bf A}_{09},$ ${\bf A}_{10},$ ${\bf A}_{11},$ ${\bf A}_{13}^{0},{\bf A}_{15},$ ${\bf A}_{17}^{0}$ and ${\bf A}_{18}^{0}$ are not rigid in the variety of $3$-dimensional standard algebras. Let us give some trivial degenerations:
${\bf A}_{02}^{\frac{1}{t}}\xrightarrow{(e_{1}+e_{2},2e_{3},-te_{1}+te_{2})}{\bf A}_{28},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1}+e_{3},e_{3},te_{2})}{\bf A}_{13}^{\alpha},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1},e_{3},te_{2})}{\bf A}_{18}^{\alpha},$  ${\bf A}_{28}\xrightarrow{(te_{2},te_{1},e_{3})}{\bf A}_{20}.$
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}({\bf A}_{04})$ $=$ $9,$
$\dim\mathcal{O}({\bf A}_{12})$ $=$ $8,$
$\dim\mathcal{O}({\bf A}_{16})=\dim\mathcal{O}({\bf A}_{17}^{\frac{1}{2}})=\dim\mathcal{O}({\bf A}_{19}^{0})=\dim\mathcal{O}({\bf A}_{19}^{\frac{1}{2}})=\dim\mathcal{O}({\bf A}_{19}^{-\frac{1}{2}})$ $=$ $7,$
$\dim\mathcal{O}({\bf A}_{02}^{\alpha})$ $=$ $6,$
$\dim\mathcal{O}({\bf A}_{14}^{0,\frac{1}{2}})=\dim\mathcal{O}({\bf A}_{14}^{0,-\frac{1}{2}})=\dim\mathcal{O}({\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}})$ $=$ $5,$
$\dim\mathcal{O}({\bf A}_{14}^{0,0})=\dim\mathcal{O}({\bf A}_{14}^{\frac{1}{2},\frac{1}{2}})=\dim\mathcal{O}({\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}})$ $=$ $3$
Algebras ${\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{16}$ and ${\bf A}_{19}^{0}$ are commutative. Hence,
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{16},$ ${\bf A}_{19}^{0}\}$$\not\to$ ${\bf A}_{19}^{\frac{1}{2}},$ ${\bf A}_{19}^{-\frac{1}{2}},$$\{{\bf A}_{17}^{\frac{1}{2}},$ ${\bf A}_{14}^{0,\frac{1}{2}},$ ${\bf A}_{14}^{0,-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},$ ${\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}},$ ${\bf A}_{02}^{\alpha}\}.$
Below we have listed all the important reasons for necessary non-degenerations.
Non-degenerations reasons
$\begin{array}[]{llll}{\bf A}_{17}^{\frac{1}{2}}\\ \end{array}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{02}^{\alpha},\ {\bf A}_{14}^{0,-\frac{1}{2}},\ {\bf A}_{14}^{0,\frac{1}{2}},\\ {\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}},\ {\bf A}_{14}^{0,0},\\ {\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},\ {\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\\ c_{22}^{1}=c_{22}^{3}=0,\ c_{21}^{3}=c_{12}^{3},\ c_{12}^{2}c_{12}^{3}=c_{11}^{3}c_{22}^{2},\\ c_{31}^{3}=c_{12}^{2}=c_{21}^{2},\ c_{13}^{3}=c_{11}^{1}\par\par\end{array}\right\}$
$\begin{array}[]{llll}{\bf A}_{19}^{\frac{1}{2}},{\bf A}_{19}^{-\frac{1}{2}}\end{array}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{02}^{\alpha},\ {\bf A}_{14}^{0,-\frac{1}{2}},\ {\bf A}_{14}^{0,\frac{1}{2}},\\ {\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}},\ {\bf A}_{14}^{0,0},\\ {\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},\ {\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\\ A_{2}A_{2}\subset A_{3},\ c_{32}^{3}=c_{23}^{3},\ c_{31}^{3}=c_{13}^{3},\\ c_{21}^{3}+c_{12}^{2}=c_{11}^{1},\ (c_{23}^{3})^{2}=c_{22}^{2}c_{33}^{3},\\ (c_{13}^{3})^{2}=c_{11}^{1}c_{13}^{3}+c_{11}^{3}c_{33}^{3},\ c_{11}^{3}c_{22}^{3}=c_{21}^{3}c_{12}^{3}\\ \end{array}\right\}$
$\begin{array}[]{llll}{\bf A}_{02}^{\alpha}\\ \end{array}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{14}^{0,0},\ {\bf A}_{14}^{\frac{1}{2},\frac{1}{2}},\\ {\bf A}_{14}^{-\frac{1}{2},-\frac{1}{2}},\ {\bf A}_{14}^{0,-\frac{1}{2}},\\ {\bf A}_{14}^{0,\frac{1}{2}},\ {\bf A}_{14}^{\frac{1}{2},-\frac{1}{2}}\end{array}$ =$\mathcal{R}=\left\{A_{1}^{3}=0\right\}$
$\begin{array}[]{llll}{\bf A}_{14}^{A,B\neq A}\end{array}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{14}^{\alpha,\alpha}\end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{2}^{2}=0,\ c_{11}^{2}=c_{11}^{3}=0,\\ (2c_{12}^{2}-c_{11}^{1}(1+2A))(2c_{12}^{2}-c_{11}^{1}(1+2B))=-4c_{13}^{2}c_{12}^{3},\\ c_{13}^{3}+c_{12}^{2}=c_{11}^{1}(1+A+B)\par\par\end{array}\right\}$
2.4. The geometric classification of noncommutative Jordan algebras
The main result of the present section is the following theorem.
Theorem 23.
The variety of complex $3$-dimensional noncommutative Jordan algebras has dimension $9$ and it has $8$ irreducible components defined by
$\mathcal{C}_{1}=\overline{\mathcal{O}({\bf A}_{04})},$  $\mathcal{C}_{2}=\overline{\mathcal{O}({\bf A}_{12})},$  $\mathcal{C}_{3}=\overline{\mathcal{O}({\bf A}_{16})},$  $\mathcal{C}_{4}=\overline{\mathcal{O}({\bf A}_{17}^{\alpha})},$
$\mathcal{C}_{5}=\overline{\mathcal{O}({\bf A}_{19}^{\alpha})},$  $\mathcal{C}_{6}=\overline{\mathcal{O}({\bf A}_{24}^{\alpha})},$  $\mathcal{C}_{7}=\overline{\mathcal{O}({\bf A}_{30})},$  $\mathcal{C}_{8}=\overline{\mathcal{O}({\bf A}_{32})}.$
In particular, there are only five rigid algebras in this variety.
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim\mathcal{O}({\bf A}_{04})=\dim\mathcal{O}({\bf A}_{24}^{\alpha})=\dim\mathcal{O}({\bf A}_{32})$ $=$ $9,$
$\dim\mathcal{O}({\bf A}_{12})=\dim\mathcal{O}({\bf A}_{17}^{\alpha})=\dim\mathcal{O}({\bf A}_{19}^{\alpha})$ $=$ $8,$
$\dim\mathcal{O}({\bf A}_{16})=\dim\mathcal{O}({\bf A}_{30})$ $=$ $7$
Below we have listed all the important reasons for necessary non-degenerations.
${\bf A}_{04}$ and ${\bf A}_{12}$ are commutative, ${\bf A}_{24}^{\alpha}$ are anticommutative, ${\bf A}_{32},$ ${\bf A}_{30},$ ${\bf A}_{19}^{\alpha}$ and ${\bf A}_{17}^{\alpha}$ are noncommutative and non-anticommutative. Hence,
$\{{\bf A}_{04},$ ${\bf A}_{12},$ ${\bf A}_{24}^{\alpha}\}$$\not\to$ $\{{\bf A}_{32},$ ${\bf A}_{30},$ ${\bf A}_{19}^{\alpha},$ ${\bf A}_{17}^{\alpha}\}.$
Non-degenerations reasons
${\bf A}_{17}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16},{\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\ A_{3}A_{3}=0,\\ c_{22}^{3}=0,\ c_{21}^{2}=c_{12}^{2},\ c_{12}^{2}c_{23}^{3}+c_{11}^{1}c_{32}^{3}=c_{13}^{3}(c_{23}^{3}+c_{32}^{3}),\\ c_{23}^{3}+c_{32}^{3}=c_{22}^{2},\ c_{21}^{3}c_{32}^{3}=c_{12}^{3}c_{23}^{3},\ c_{12}^{2}c_{21}^{3}=c_{11}^{3}c_{23}^{3}\par\par\end{array}\right\}$
${\bf A}_{19}^{\alpha}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{16},{\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}A_{1}A_{2}+A_{2}A_{1}\subset A_{2},\ A_{1}A_{3}+A_{3}A_{1}\subset A_{3},\\ \ A_{2}A_{3}+A_{3}A_{2}=0,\ c_{31}^{3}+c_{13}^{3}=c_{11}^{1}\end{array}\right\}$
${\bf A}_{32}$$\not\rightarrow$ $\begin{array}[]{llll}{\bf A}_{12},\ {\bf A}_{16},\ {\bf A}_{17}^{\alpha},\\ {\bf A}_{19}^{\alpha},\ {\bf A}_{30}\\ \end{array}$ =$\mathcal{R}=\left\{\begin{array}[]{lllll}c_{11}^{2}=0,\ c_{11}^{3}=0,\ c_{22}^{1}=0,\ c_{22}^{3}=0,\ c_{33}^{1}=0,\ c_{33}^{2}=0,\\ c_{12}^{3}=-c_{21}^{3},\ c_{21}^{1}+c_{12}^{1}=c_{22}^{2},\ c_{21}^{2}+c_{12}^{2}=c_{11}^{1},\\ c_{23}^{3}+c_{32}^{3}=c_{22}^{2},\ c_{31}^{3}+c_{13}^{3}=c_{11}^{1}\par\par\par\par\end{array}\right\}$
All necessary degenerations are given below:
  1. (1)
    ${\bf A}_{19}^{\frac{\alpha}{2}}\xrightarrow{(te_{1},-2t^{2}e_{2}+e_{3},t^{3}e_{2})}{\bf A}_{02}^{\alpha},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1}+e_{2},e_{3},te_{2})}{\bf A}_{13}^{\alpha},$  ${\bf A}_{33}^{\beta}\xrightarrow{(e_{1}+(\beta-\alpha)e_{2}-\alpha e_{3},t^{2}e_{3},te_{2}+te_{3})}{\bf A}_{14}^{\alpha,\beta},$  ${\bf A}_{19}^{\alpha}\xrightarrow{(e_{1},e_{3},te_{2})}{\bf A}_{18}^{\alpha},$  ${\bf A}_{17}^{\frac{-2-t}{2t}}\xrightarrow{(e_{1}+e_{2},te_{2},e_{3})}{\bf A}_{29},$  ${\bf A}_{33}^{\frac{2\alpha-t}{2+2t}}\xrightarrow{((1+t)e_{1}-(\alpha+\frac{t}{2})e_{3},-t^{2}e_{3},e_{2}+te_{3})}{\bf A}_{31}^{\alpha},$  ${\bf A}_{32}\xrightarrow{(e_{1}-\alpha e_{3},te_{3},(\alpha+t^{2})e_{2}+e_{3})}{\bf A}_{33}^{\alpha}.$
  2. (2)
    Thanks to Theorem 20, all irreducible components of the variety of $3$-dimensional Kokoris algebras are defined by the following algebras ${\bf A}_{02}^{\alpha},$  ${\bf A}_{04},$  ${\bf A}_{24}^{\alpha},$  ${\bf A}_{29},$ and ${\bf A}_{30}.$ Hence, ${\bf A}_{01},$ ${\bf A}_{03},$ ${\bf A}_{05},$ ${\bf A}_{06},$ ${\bf A}_{07},$ ${\bf A}_{08},$ ${\bf A}_{09},$ ${\bf A}_{10},$ ${\bf A}_{11},$ ${\bf A}_{20},$ ${\bf A}_{21},$ ${\bf A}_{22}^{\alpha},$ ${\bf A}_{23},$ ${\bf A}_{25},$ ${\bf A}_{26},$ ${\bf A}_{27}$ and ${\bf A}_{28}$ do not give irreducible components in the variety of noncommutative Jordan algebras.
  3. (3)
    Thanks to [12], all irreducible components the variety of $3$-dimensional Jordan algebras are defined by ${\bf A}_{04},{\bf A}_{12},{\bf A}_{14}^{0,0},{\bf A}_{16},{\bf A}_{19}^{0}.$ Hence, ${\bf A}_{15}$ does not give an irreducible component in the variety of noncommutative Jordan algebras.